rodrilr7268:

what is 0 divivded by 0

Hoodmemes:

Its (0) simple math

rodrilr7268:

-0

Hoodmemes:

Woah,woah u fr?

KazGotFetee:

0 / 0 is 0

Hoodmemes:

Its "undefined" there is no answer

KazGotFetee:

So 0 / 0 =(?+?)

KazGotFetee:

or = Undefined

KazGotFetee:

lol

KazGotFetee:

Dis considered spam?

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @KazGotFetee So 0 / 0 =(?+?) \(\color{#0cbb34}{\text{End of Quote}}\) like look The unfunny joke on the qc cal and then the google one

KazGotFetee:

... holy....

jhonyy9:

\[\frac{ 0 }{ 0 }\] hope you know that a fraction with denominator equal zero is always undefined @AZ opinion pls. ? ty